OpenStudy (anonymous):

how many times does 1/2 go in to 2/3

OpenStudy (anonymous):

1

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

just divide 1/2 = 0.5 2/3 = 0.666666... 0.5 goes into 0.6 one time

OpenStudy (anonymous):

im not sure thats correct

OpenStudy (anonymous):

Why not?

OpenStudy (anonymous):

1/2 = %50 2/3 = %66.66666... 1/2 only goes into 2/3 once